# On the Analysis of Misspecified Variational Inequalities with Nonlinear Constraints

### v1

@@ 5 Numerical Experiment
$C^{\prime}_{T}\triangleq\Big(\tfrac{1}{2\gamma}+\tfrac{3}{2}L_{F,x}\Big)D_{X}^{2}+2L_{F,\theta}D_{X}D_{\Theta}\quad\text{such that}\quad\sum_{k=0}^{K-1}\mathcal{T}_{k}(x)\leq C^{\prime}_{T}<\infty\quad\forall K,\forall x\in X.$ (35)
For $\mathcal{R}_{k}(x,\lambda)$, using $\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1})|\leq JD_{f}(\theta^{*})$, together with the assumptions stated in Lemma 4.3 and its result, we define the uniform bound $\|\lambda_{k}-\lambda^{*}\|\leq D_{\Lambda}$. Hence, we obtain
$C^{\prime}_{R}(x,\lambda)$ $\triangleq\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+L^{\Phi}_{\lambda\theta}(D_{\Lambda}+\|\lambda-\lambda^{*}\|)+\rho L_{f,\theta}\,JD_{f}(\theta^{*})\Big)\sum_{k}\|\theta_{k}-\theta^{*}\|$
$\quad+2L_{F,\theta}\sum_{k}\|\theta_{k}-\theta^{*}\|^{2},$ (36)
satisfying $\sum_{k=0}^{K-1}\mathcal{R}_{k}(x,\lambda)\leq C^{\prime}_{R}(x,\lambda)\quad\text{for all }k.$ Using (34), (35) and (D) in (33), we get
@@
Apply (D) with $x=x^{*}$. Since $x^{*}\in\mathcal{X}(\theta^{*})$, we have $\Phi_{\rho}(x^{*},\lambda_{k},\theta^{*})\leq 0$ for all $k$, hence
@@
Choose $\tilde{\lambda}_{j}\triangleq\begin{cases}1+\lambda_{j}^{*},&\text{if }f_{j}(\bar{x}_{K},\theta^{*})>0\\ 0,&\text{otherwise}\end{cases}$, so that $f(\bar{x}_{K},\theta^{*})^{\top}\tilde{\lambda}-[f(\bar{x}_{K},\theta^{*})]_{+}^{\top}\lambda^{*}=\mathbf{1}^{\top}[f(\bar{x}_{K},\theta^{*})]_{+}$. Using also the boundedness of $\{\lambda_{k}\}$, in (40) with $\lambda=\tilde{\lambda}$ gives
$\mathbf{1}^{\top}[f(\bar{x}_{K},\theta^{*})]_{+}\leq\frac{C^{\prime}_{T}+C^{\prime}_{R}(x^{*},\tilde{\lambda})}{K}+\frac{(\sqrt{J}+2D_{\Lambda})^{2}}{2\rho K}\triangleq\frac{C_{\mathrm{feas}}}{K},$ (41)
where $C_{\mathrm{feas}}=C^{\prime}_{T}+\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+L^{\Phi}_{\lambda\theta}(\sqrt{J}+2D_{\Lambda})+\rho L_{f,\theta}JD_{f}(\theta^{*})\Big)\sum_{k}\|\theta_{k}-\theta^{*}\|+2L_{F,\theta}\sum_{k}\|\theta_{k}-\theta^{*}\|^{2}+\frac{(\sqrt{J}+2D_{\Lambda})^{2}}{2\rho}$. Now if we define $\epsilon\triangleq\frac{C_{\mathrm{feas}}}{K},$ and the enlarged feasible set $\mathcal{X}_{\epsilon}(\theta^{*})=\{x\in X\mid\mathbf{1}^{\top}[f(x,\theta^{*})]_{+}\leq\epsilon\}.$ Then $\sum_{j=1}^{J}[f_{j}(x,\theta^{*})]_{+}\leq\epsilon$, and using the definition of $\Phi_{\rho}$ together with $\|\lambda_{k}\|\leq D_{\Lambda}$ yields
@@
Note that, this bound is uniform in $k$. Therefore, we have
$\frac{1}{K}\sum_{k=0}^{K-1}\Phi_{\rho}(x,\lambda_{k},\theta^{*})\leq D_{\Lambda}\epsilon+\frac{\rho}{2}\epsilon^{2}$
Applying (D) with this $x$ and $\lambda=0$ yields
@@
$\leq\frac{C^{\prime}_{T}+C^{\prime}_{R}(x,0)}{K}+D_{\Lambda}\epsilon+\frac{\rho}{2}\epsilon^{2}$
$\leq\frac{C^{\prime}_{T}+C^{\prime}_{R}(x,0)}{K}+D_{\Lambda}\frac{C_{\mathrm{feas}}}{K}+\frac{\rho}{2}\left(\frac{C_{\mathrm{feas}}}{K}\right)^{2}$ (42)
@@
where $C^{\prime}_{T}$, and $C_{\mathrm{feas}}$ are defined as above and $C^{\prime}_{R}(x,0)=\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+2L^{\Phi}_{\lambda\theta}D_{\Lambda}+\rho L_{f,\theta}JD_{f}(\theta^{*})\Big)\sum_{k}\|\theta_{k}-\theta^{*}\|+2L_{F,\theta}\sum_{k}\|\theta_{k}-\theta^{*}\|^{2}.$

### v3

@@ 5 Numerical Experiment
$\sum_{k=0}^{K-1}\mathcal{T}_{k}(x)\leq\tfrac{1}{2\gamma}D_{X}^{2}+2L_{F,\theta}D_{X}D_{\Theta}\triangleq C^{\prime}_{T}\quad\forall K,\quad\forall x\in X.$ (43)
Next, we focus on $\sum_{k=0}^{K-1}\mathcal{R}_{k}(x,\lambda)$. By continuity on the compact sets $X,\Theta$ we have that $\sup_{x\in X,\theta\in\Theta}|f_{j}(x,\theta^{*})|\leq D_{f}$ for all $j=1,\cdots,J$. Now, let the constant-parameters $\rho_{k}\equiv\rho>0$ and $\gamma_{k}\equiv\gamma>0$. By Lemma 4.4, the dual sequence is bounded, and hence there exists $D_{\Lambda}>0$ such that $\|\lambda_{k}-\lambda^{*}\|\leq D_{\Lambda}$ for all $k\geq 0.$ Therefore, $\frac{C_{2}}{2}\|\lambda_{k}-\lambda^{*}\|\,\|x_{k+1}-x_{k}\|^{2}\leq\frac{C_{2}D_{\Lambda}}{2}\|x_{k+1}-x_{k}\|^{2}$. Consequently, if the primal step-size is chosen so that $\gamma\in\Bigl(0,\,\bigl(\rho C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}+C_{2}D_{\Lambda}\bigr)^{-1}\Bigr),$ then the whole coefficient of $\|x_{k+1}-x_{k}\|^{2}$ in $\mathcal{R}_{k}(x,\lambda)$ is non-positive and may be discarded. The remaining terms in $\mathcal{R}_{k}(x,\lambda)$ are controlled by the boundedness of $X$, $\Theta$, and $\{\lambda_{k}\}_{k\geq 0}$ together with the learning-error summability assumption. In particular, since $\|\lambda_{k}\|\leq\|\lambda_{k}-\lambda^{*}\|+\|\lambda^{*}\|\leq D_{\Lambda}+\|\lambda^{*}\|$, the last term $2L_{f,\theta}(J\rho D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|$ is bounded by $2L_{f,\theta}\bigl(J\rho D_{f}+\sqrt{J}(D_{\Lambda}+\|\lambda^{*}\|)\bigr)\|\theta_{k}-\theta^{*}\|$ which together with $\frac{J\rho}{2}L_{f,\theta}^{2}\|\theta_{k}-\theta^{*}\|^{2}$ are summable by the learning-error assumption. Therefore, we conclude that $\sum_{k=0}^{K-1}\mathcal{R}_{k}(x,\lambda)\leq C^{\prime}_{R}(x,\lambda)$ where
$C^{\prime}_{R}(x,\lambda)\triangleq$ $\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+2L^{\Phi}_{\lambda\theta}D_{\Lambda}+2L_{f,\theta}\big(J\rho D_{f}+\sqrt{J}(D_{\Lambda}+\|\lambda^{*}\|)\big)\Big)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|$
$\quad+\left(L_{F,\theta}+\frac{J\rho}{2}L_{f,\theta}^{2}\right)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}<+\infty.$ (44)
Using (43) and (E) in (E), we get
@@
Let $x=x^{*}$ within (E). Since $x^{*}\in\mathcal{X}(\theta^{*})$, we have $\Phi_{\rho}(x^{*},\lambda_{k},\theta^{*})\leq 0$ for all $k$, hence
@@
Let $\lambda=\tilde{\lambda}$ within (48) such that $\tilde{\lambda}_{j}\triangleq\begin{cases}1+\lambda_{j}^{*},&\text{if }f_{j}(\bar{x}_{K},\theta^{*})>0\\ 0,&\text{otherwise}\end{cases}$, hence, $f(\bar{x}_{K},\theta^{*})^{\top}\tilde{\lambda}-[f(\bar{x}_{K},\theta^{*})]_{+}^{\top}\lambda^{*}=\mathbf{1}^{\top}[f(\bar{x}_{K},\theta^{*})]_{+}$ which implies that
$\mathbf{1}^{\top}[f(\bar{x}_{K},\theta^{*})]_{+}\leq\frac{C^{\prime}_{T}+C^{\prime}_{R}(x^{*},\tilde{\lambda})}{K}+\frac{J+\|\lambda_{0}-\lambda^{*}\|^{2}}{\rho K}=\frac{C_{\mathrm{feas}}}{K},$ (49)
where $C_{\mathrm{feas}}=C^{\prime}_{T}+\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+2L^{\Phi}_{\lambda\theta}D_{\Lambda}+2L_{f,\theta}\big(J\rho D_{f}+\sqrt{J}(D_{\Lambda}+\|\lambda^{*}\|)\big)\Big)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|+\left(L_{F,\theta}+\frac{J\rho}{2}L_{f,\theta}^{2}\right)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}+\frac{J+\|\lambda_{0}-\lambda^{*}\|^{2}}{\rho}$.
Moreover, if we define $\epsilon\triangleq\frac{C_{\mathrm{feas}}}{K}$ and the enlarged feasible set $\mathcal{X}_{\epsilon}(\theta^{*})=\{x\in X\mid\mathbf{1}^{\top}[f(x,\theta^{*})]_{+}\leq\epsilon\}$. Then, $\sum_{j=1}^{J}[f_{j}(x,\theta^{*})]_{+}\leq\epsilon$, and using the definition of $\Phi_{\rho}$ together with $\|\lambda_{k}\|\leq D_{\Lambda}$ yields
@@
Note that this bound is uniform in $k$. Therefore, we have $\frac{1}{K}\sum_{k=0}^{K-1}\Phi_{\rho}(x,\lambda_{k},\theta^{*})\leq D_{\Lambda}\epsilon+\frac{\rho}{2}\epsilon^{2}.$ Now, let $\lambda=\mathbf{0}$ in (E) to conclude that
@@
$\leq D_{\Lambda}\epsilon+\frac{\rho}{2}\epsilon^{2}+\frac{C^{\prime}_{T}+C^{\prime}_{R}(x,0)}{K}$
$\leq D_{\Lambda}\frac{C_{\mathrm{feas}}}{K}+\frac{\rho}{2}\left(\frac{C_{\mathrm{feas}}}{K}\right)^{2}+\frac{C^{\prime}_{T}+C^{\prime}_{R}(x,0)}{K}.$ (50)
@@
where $C^{\prime}_{T}$, and $C_{\mathrm{feas}}$ are defined as above and $C^{\prime}_{R}(x,0)=\Big(L_{F,\theta}D_{X}+2L^{\Phi}_{x\theta}D_{X}+2L^{\Phi}_{\lambda\theta}D_{\Lambda}+2L_{f,\theta}\big(J\rho D_{f}+\sqrt{J}(D_{\Lambda}+\|\lambda^{*}\|)\big)\Big)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|+\left(L_{F,\theta}+\frac{J\rho}{2}L_{f,\theta}^{2}\right)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}.$